### v1

# GraphRAG-R1: Graph Retrieval-Augmented Generation with Process-Constrained Reinforcement Learning

Abstract.
Graph Retrieval-Augmented Generation (GraphRAG) has shown great effectiveness in enhancing the reasoning abilities of Large Language Models (LLMs) by leveraging graph structures for knowledge representation and modeling complex real-world relationships. However, existing GraphRAG methods still face significant bottlenecks when handling complex problems that require multi-hop reasoning, as their query and retrieval phases are largely based on pre-defined heuristics and do not fully utilize the reasoning potentials of LLMs. To address this problem, we propose GraphRAG-R1, an adaptive GraphRAG framework by training LLMs with process-constrained outcome-based reinforcement learning (RL) to enhance the multi-hop reasoning ability. Our method can decompose complex problems, autonomously invoke retrieval tools to acquire necessary information, and perform effective reasoning. Specifically, we utilize a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking capability to train the model. Next, we design two process-constrained reward functions. To handle the shallow retrieval problem, we design a Progressive Retrieval Attenuation (PRA) reward to encourage essential retrievals. Then, to handle the over-thinking problem, we design Cost-Aware F1 (CAF) reward to balance the model performance with computational costs. We further design a phase-dependent training strategy, containing three training stages corresponding to cold start and these two rewards. These stages empower GraphRAG with format following, behavior shaping, and smartness optimization abilities, respectively. Lastly, our method adopts a hybrid graph-textual retrieval to improve the reasoning capacity. Extensive experimental results demonstrate that GraphRAG-R1 significantly boosts LLM capabilities in solving complex reasoning problems compared to state-of-the-art GraphRAG methods on both in-domain and out-of-domain datasets. Furthermore, our framework can be flexibly integrated with various existing retrieval methods, consistently delivering performance improvements.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Large Language Models (LLMs) have attracted ever-growing attention in the past few years due to their outstanding ability in answering complex questions and generating human-quality text (Vaswani et al., 2017; WANG et al., 2024). Despite remarkable progress, LLMs still exhibit significant limitations when faced with complex tasks that require integrating vast and diverse knowledge domains or require deep reasoning and multi-step logical inference (Talmor et al., 2019; Petroni et al., 2021). As shown in Figure 1(a), LLMs will provide an incorrect response when the question is over-complex and requires knowledge beyond their boundary. Graph Retrieval-Augmented Generation (GraphRAG) (Xiang et al., 2025; Zhang et al., 2025a; Peng et al., 2024; Han et al., 2024; Zhu et al., 2024; Li et al., 2025), which utilizes knowledge graphs for knowledge representation, enables effective modeling of complex real-world relationships, such as causality, hierarchy, and dependency (Edge et al., 2024; Peng et al., 2024).
However, existing GraphRAG methods still face significant performance bottlenecks when handling complex problems that require multi-hop reasoning. The main bottleneck is that existing GraphRAG methods rely on pre-defined or heuristic rules during query processing and graph retrieval, which are too rigid and inflexible. Concretely, during query processing, the semantic or lexical similarity search (Han et al., 2024; Gao et al., 2023b; Fan et al., 2024) remains the most widely adopted. For semantic similarity search, queries with high complexity are averaged during vectorization, making it difficult to gather accurate information for each distinct semantic element. On the other hand, lexical similarity search captures only the superficial meaning of the query, easily losing implicit relationships and reasoning logic. Therefore, as shown in Figure 1(b), existing GraphRAG methods are unable to answer the query, since the retrieval results are incomplete or incorrect when queries are overly complex. Besides, during the graph retrieval phase, current methods employ a combination of graph traversal and subgraph retrieval for path reasoning, which are rather rigid (Sun et al., [n. d.]; Ma et al., 2024; Jimenez Gutierrez et al., 2024; Guo et al., 2024; Li et al., 2024a). For instance, ToG (Sun et al., [n. d.]; Ma et al., 2024), a representative GraphRAG method, starts from a node in the knowledge graph and traverses the surrounding nodes based on breadth-first search strategy with pruning. While this strategy might work for short paths, it struggles with long-path problems, which require capturing logically sound but distant or even disconnected relationships. This inflexibility often leads to broken reasoning chains and failed retrievals, making it unsuitable for multi-hop problems with complex reasoning logic. As a result, the existing GraphRAG methods fail to fully leverage the LLM’s deep reasoning potential.
In this paper, we try to further enhance the multi-hop reasoning ability of GraphRAG. Reinforcement learning (RL) has recently been shown to be effective in endowing LLMs with powerful reasoning abilities. For example, in the case of DeepSeek-R1 (Guo et al., 2025), it has been discovered that large-scale reinforcement learning can endow models with deep thinking capabilities (Xie et al., 2025; Jaech et al., 2024; Team et al., 2025), even without any supervised fine-tuning. This effectiveness may result from RL’s use of reward signals to encourage exploration, refining dynamic reasoning strategies from static knowledge (Ouyang et al., 2022) beyond static reasoning patterns (Asai et al., 2024; Zhang et al., 2025b; Xiong et al., 2025; Guan et al., 2025). However, how to utilize RL for GraphRAG training remains unexplored in the literature. The major technical obstacle lies in how to design proper reward functions so that the reasoning process of LLMs within GraphRAG can be properly constrained, avoiding potential problems such as shallow retrieval (Helwe et al., 2021; Deng et al., 2025) and over-thinking (Sui et al., 2025; Chen et al., 2024; Cuesta-Ramirez et al., 2025).
To address this issue, we propose GraphRAG-R1, an adaptive GraphRAG framework with deep reasoning capabilities through process-constrained outcome-based RL. As shown in Figure 1(c), our method can decompose complex multi-hop problems by reasoning and autonomous, on-demand invocation of retrieval tools. Specifically, we employ a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking mechanism (Song et al., 2025) to autonomously leverage external retrieval tools on demand. Next, we design two process-constrained reward functions, Progressive Retrieval Attenuation (PRA) and Cost-Aware F1 (CAF), aiming to inject prior knowledge into the LLM’s training process. In detail, PRA aims to dynamically adjust the reward for retrieval actions. Initially, it encourages necessary retrievals to gather key information, preventing the model from shallow retrieval. As the reasoning deepens, PRA progressively attenuates retrieval rewards, guiding the model to rely more on acquired information for internal reasoning and avoid unnecessary retrieval overhead. On the other hand, CAF evaluates the final answer correctness using the F1 score while also accounting for the computational resources consumed during the reasoning and retrieval process. It encourages the model to achieve correct results more efficiently to prevent over-thinking. These two rewards collectively balance between effectiveness and computational cost. We further design a three-stage phase-dependent training strategy. In the first stage, we focus on the answer format to alleviate cold start problems. In the second and third stage, we focus on improving the answer quality using the aforementioned two rewards. Lastly, our method employs a hybrid graph-textual retrieval instead of only retrieving textual information as in existing GraphRAGs. By leveraging graph structures to precisely capture entity relationships while preserving the rich semantic details of the original text, our method provides more comprehensive and integrated knowledge for complex reasoning.
Experiments show that GraphRAG-R1 significantly outperforms existing GraphRAG methods in solving complex reasoning problems, across both in-domain and out-of-domain datasets. Additionally, our framework offers flexible integration with various retrieval methods, consistently enhancing their performance.
Our main contributions are summarized as follows:
  • We propose GraphRAG-R1, a pioneering framework that utilizes GRPO that supports rollout-with-thinking to enhance the reasoning ability of GraphRAG. To the best of our knowledge, we are the first to explore RL-based LLM training in GraphRAG.
  • We propose PRA and CAF, two tailored reward designs to guide the training process of RL and handle the shallow retrieval and over-thinking problems. We also propose a phase-dependent training strategy and a hybrid graph-textual retrieval.
  • Experiments demonstrate that our method exhibits state-of-the-art performance on both in-domain and out-of-domain datasets, and can be flexibly integrated with various existing retrieval methods with consistently improved performance.
Figure 1. An example of comparing using only LLMs, GraphRAG, and GraphRAG-R1 in answering complex problems: (a) The LLM (Qwen-2.5-7b) directly produces an output, but the answer is incorrect, (b) HippoRAG2 (Gutiérrez et al., 2025) enhances the LLM (Qwen-2.5-7b) by retrieving external knowledge, but fail to generate correct outputs for complex problem, (c) our method successfully decomposes the problem and autonomously invokes retrieval methods, producing the correct output.
2. Related Work
2.1. Graph Retrieval-Augmented Generation
Graph Retrieval-Augmented Generation (GraphRAG) offers a novel paradigm to enhance the reasoning abilities of LLMs by explicitly modeling entity relationships through graph structure such as knowledge graphs (Peng et al., 2024; Han et al., 2024). GraphRAG has been widely adopted across multiple domains. For instance, GraphRAG based on knowledge graph (Pan et al., 2024) incorporates techniques like subgraph retrieval and Graph Neural Networks (GNNs) (He et al., 2024) to enhance the accuracy and explainability of LLMs in complex knowledge reasoning question answering. GraphRAG based on document graphs (Sonawane and Kulkarni, 2014) model semantic relationships between documents and internal document structures, effectively supporting long-document summarization and cross-document information integration (Sarthi et al., [n. d.]). Furthermore, GraphRAG for specialized domains are increasingly trending. Examples include leveraging social graphs (Jiang and Ferrara, 2023; Zeng et al., 2024) to analyze user relationship for precision recommendations and constructing medical or disease graphs (Yang et al., 2024b; Li et al., 2024b) that integrate multi-source heterogeneous data to enable personalized treatment plan generation.
However, existing GraphRAG methods still face significant performance bottlenecks when facing complex reasoning problem, primarily due to pre-defined or heuristic rules during query processing and graph retrieval that struggle with highly complex multi-hop problems. For example, ToG (Sun et al., [n. d.]; Ma et al., 2024) introduces reasoning steps using breadth-first search (BFS) with pruning approach, but it heavily relies on prompt engineering and external LLMs and thus lacks flexibility. G-Retriever (He et al., 2024) adopts K-Nearest Neighbors-based retrieval with limited efficacy in multi-hop scenarios. Moreover, existing approaches suffer from shallow retrieval depth, since they often depend on predefined fixed paths or hop counts. HippoRAG (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025) optimizes knowledge integration through personalized PageRank but and only supports single-hop retrieval. KGP (Wang et al., 2024) and LightRAG (Guo et al., 2024) are typically confined to 2-3 hops, exhibiting significant performance degradation when handling problems that require deep-dependence mining. In contrast, our work improves the reasoning ability of GraphRAG for complex problems using process-constrained RL, aiming to effectively resolves the issues of rigid retrieval strategies and complex reasoning shortcomings in existing GraphRAG approaches.
2.2. RL-Enhanced Retrieval Scheme
Recently, reinforcement learning (RL) has become increasingly important for complex reasoning (Xie et al., 2025; Yang et al., 2024a; Ouyang et al., 2022). For example, DeepSeek-R1 (Guo et al., 2025) shows that large-scale RL can endow LLMs with deep thinking capabilities. Some pioneering work have also adopted RL to enhance RAG (Jin et al., 2025; Song et al., 2025). For example, R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) achieve co-training of retrieval and generation, effectively improving the model’s ability to solve complex problems. However, a systematic exploration of RL schemes within the GraphRAG domain remains open. Technically speaking, directly extending existing RL methods into GraphRAG faces several challenges, Firstly, the existing methods usually employ a single outcome-oriented reward function, e.g., based solely on the correctness of final answer. However, this design becomes too simple for GraphRAG, since complex problem may involve multiple different retrieval calls to different components of the external knowledge graph. More critically, existing methods have often encountered with reward hacking in RL training, potentially resulting in shallow retrieval and over-thinking in GraphRAG. To tackle these issues, we propose GraphRAG-R1 with tailored reward functions and training strategy to non-trivially explore the potential of RL in enhancing GraphRAG, to the best of our knowledge, for the first time.
3. Method
In this section, we present our proposed method. First, we introduce the rollout retrieval enhanced GRPO algorithm. Second, we detail the designs of process-constrained reward , which includes PRA and CAF, together with the phase-dependent training strategy. Finally, we present our hybrid graph-textual retrieval design. An overall framework of our method is shown in Figure 2.
Figure 2. An overview of GraphRAG-R1 : (a) the rollout retrieval enhanced GRPO as the training strategy of LLM, (b) process-constrained reward designs, containing PRA and CAF rewards and the phase-dependent training strategy, (c) the hybrid graph-textual retrieval, which is more informative than text fragments.
3.1. Rollout Retrieval Enhanced GRPO
In this subsection, we introduce the details of our training method, which is a modified version of GRPO that supports the roll-out-with-thinking capability (Song et al., 2025).
Our goal is to optimize the LLM using RL to generate high-quality outputs augmented by an external retrieval tool $\mathcal{R}$ that assists the reasoning generation process. Specifically, denote the policy of RL as $\pi_{\theta}(\cdot)$ parametrized by $\theta$. Let $q$ represent input queries drawn from the training dataset $\mathcal{D}$, $o$ represent the generated outputs, and $r(\cdot)$ be the reward function. We formalize the objective as follows:
(1) $\max_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D},o\sim\pi_{\theta}(\cdot\mid q;\mathcal{R})}\big{[}r(q,o)-\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\big{]},$
where $\pi_{\text{ref}}(\cdot)$ is the reference LLM, $\mathbb{D}_{\text{KL}}(\cdot)$ is the KL-divergence measure, and $\beta$ is a hyperparameter that controls the strength of the KL-divergence regularization. More specifically, our method builds on the GRPO algorithm (Guo et al., 2025), which uses the average reward of multiple sampled outputs as a baseline. Specifically, for each input question $q$, GRPO samples a group of $G$ responses ${o_{1},o_{2},...,o_{G}}$ using the old policy $\pi_{\theta_{\text{old}}}$. The policy model is then optimized by maximizing the following objective function:
(2) $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}q\sim D\\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}\end{subarray}}\bigg{[}\frac{1}{G}\sum_{i=1}^{G}\min\bigg{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}\cdot A_{i},$
$\text{clip}\left(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\right)\cdot A_{i}\bigg{)}-\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg{]},$
where the group size $G$ determines the number of responses generated per query, $\theta_{\text{old}}$ denotes the previous policy parameters, $\epsilon$ is the clipping range hyperparameter to constrain policy updates, and $A_{i}$ represents the advantage, computed based on the relative rewards of outputs within each group as follow:
(3) $A_{i}=\frac{r_{i}-\mu_{r}}{\sigma_{r}},$
where $r_{i}$ is the reward for individual output, $\mu_{r}$ means the average reward of all $G$ outputs for the same query $q$, and $\sigma_{r}$ is the group reward standard deviation.
To better adapt GRPO to our GraphRAG scenario, we adopt two key modifications proposed by R1-Searcher (Song et al., 2025): Rollout-with-Thinking and retrieval-masked loss. The core objective is to enable the model to autonomously leverage external retrieval tools on demand during inference, i.e., accessing supplementary knowledge when needed to solve problems beyond its internal capacity, and receive rewards for generating correct answers.
3.1.1. Rollout-with-Thinking
The model is guided to utilize the external retrieval tool during generation by using two special tokens <begin_of_query>...<end_of_query> to indicate the call to the search tool. When <|end_of_query|> appears, it signifies that the retrieval query has been fully generated. At this point, model generation is paused, the external retrieval tool is called, and the retrieved content is marked with <|begin_of_documents|> ... <|end_of_documents|>. This content then guides the model’s next step of reasoning. This method ensures retrieval is seamlessly integrated into the reasoning process, allowing the model to continue its reasoning based on the retrieved documents without interruption. Detailed prompts are provided in Appendix A.4
3.1.2. Retrieval-Masked Loss
In the original GRPO, token-level loss is computed throughout the rollout sequence. In the GraphRAG scenario, the rollout sequence consists of tokens generated by the LLM and tokens retrieved from external channels. Applying the same optimization to retrieved tokens and LLM-generated tokens may negatively impact the training process. Therefore, <begin_of_documents> and <end_of_documents> are designated as special tokens and their enclosed content is masked during training, ensuring the policy gradient objective is calculated exclusively on LLM-generated tokens. This approach stabilizes training while preserving the flexibility of retrieval-augmented generation.
3.2. Process-Constrained Reward Design
In this subsection, we introduce our reward functions to constrain the retrieval process, enabling effective and efficient reasoning.
RL methods have recently attracted considerable attention to improve the reasoning abilities of LLMs. However, if simple reward functions are employed, such as relying solely on final answer correctness, existing methods easily face a technical problem known as “Reward Hacking” (Gao et al., 2023a), i.e., the model exploits certain loopholes in the reward to gain high scores rather than genuinely improving reasoning capabilities. To address this potential issue, our reward design injects controlled human prior knowledge while preserving the fundamental reward framework. Concretely, we introduce the following three rewards: format reward, PRA, and CAF, as follows.
3.2.1. Format reward
This reward function is designed to steer the model toward strictly adhering to the specified output format, ensuring that it learns to correctly invoke external retrieval and generate the final answer. The model should place the final response entirely within the <answer>...</answer> tags and guarantee that the entire output is fully readable and free of uninterpretable artifacts. When invoking retrieval, the model should encapsulate retrieval queries within the <|begin_of_query|>...<|end_of_query|> tags with retrieved information contained within <|begin_of_documents|> ... <|end_of_documents|>. The model should explicitly invoke the retrieval mechanism through this query process rather than directly generating content without retrieval. An illustration of the format is shown in Figure 4 in Appendix A.4.
The format reward is defined as follows:
(4) $R_{\text{format}}=\begin{cases}0.5,&\text{correct}\\ 0,&\text{incorrect}.\end{cases}$
3.2.2. Progressive Retrieval Attenuation (PRA)
The most straightforward design to encourage GraphRAG models to invoke retrieval calls is to adopt a single retrieval reward. However, this strategy tends to cause the model to invoke the external retrieval only once and cause the shallow retrieve problem, as some complex multi-hop problems require multiple retrieval calls. On the other hand, one can sum up rewards for multiple retrieves, but it may trigger unnecessary retrieval calls. To circumvent these issues, we propose Progressive Retrieval Attenuation (PRA), specifically designed to provide tailored rewards for each invocation while incorporating a decay factor. With each additional retrieval call, the reward undergoes exponential decay relative to the base reward, concurrently accumulating rewards from the preceding invocations. Formally, the retrieval reward is defined as follows:
(5) $R_{n}=\begin{cases}R_{0},&n=1\\ R_{n-1}+R_{0}\times k^{n-1},&n>1\end{cases},$
[…]
(6) $R_{0}\leq R_{n}<\frac{1}{1-k}R_{0},\;\forall n.$
When $k=0$, PRA degenerates into a single retrieval reward. When $k=1$, it degenerates into the sum of rewards. The final reward is $R_{\text{retrieval}}=R_{N}$, where $N$ is the actual number of retrieval calls.
Through this reward function, we can effectively restrict the number of model-triggered retrieval operations, effectively preventing shallow retrieval. For different data characteristics, the hyperparameter $k$ can be flexibly adjusted to better steer the model’s learning process.
3.2.3. Cost-Aware F1 (CAF)
Although the PRA reward can regulate the retrieval frequency of the model within a reasonable range, this mechanism alone still relies on prior knowledge and could not fully activate the reasoning capabilities of LLMs. For example, the LLM may adopt unnecessarily numerous of retrievals, leading to the “over-thinking” problem and greatly waste the computational resources. An ideal GraphRAG model should deliver high-quality responses with precisely needed retrieval operations, which requires three core competencies: accurately assessing the problem difficulty, intelligently decomposing complex queries, and dynamically optimizing retrieval strategies.
To this end, we design the CAF reward, which achieves the above objective through another decay coefficient. Formally, the CAF reward is defined as:
(7) $R_{\text{CAF}}=F1\times a\times e^{-b\times N},$
where $F1$ is the F1 score of the generated answers, $a$ and $b$ are hyperparameters to control the coefficients in the regulatory function, and $N$ denotes the total number of retrieval operations. The CAF rewards ensure that responses attaining identical F1 scores receive higher rewards when using fewer retrievals, thereby incentivizing the model to proactively recognize problem complexity and refine retrieval decisions. Consequently, CAF reward establishes a balance among reasoning depth, retrieval cost, and generation quality, effectively mitigating suboptimal behaviors of over-thinking.
3.2.4. Phase-Dependent Training Strategy
Despite the effectiveness of these reward functions, simultaneously optimizing them induces technical challenges. To address this challenge, we propose a phase-dependent training strategy to avoid the potential conflicts between objectives. Concretely, we split the training into three sequential stages, detailed as follows.
Specifically, in the first stage, we utilize another LLM to generate data following our desired format. Then, we utilize supervised fine-tuning (SFT) to train the LLM with the generated data. This training process can enable the model to follow our desired format by addressing the cold start problem and stabilizing the training process, laying the foundation for the following training process.
In the second stage, we aim to shape the behavior of the model, synergistically combining the format reward and the PRA reward. This design recognizes standardized tool-calling protocols and retrieval behavior as essential pre-requisites for high-quality reasoning. By learning with two objectives, the model can efficiently learn how and when to invoke the retrieval tools with structured output patterns. The decay mechanism of PRA dynamically regulates retrieval behavior through a high initial reward, driving essential information retrieval, while its exponential decay characteristic suppresses redundant operations.
In the last stage, we further optimize the overall smartness of the model, adopting CAF as the primary reward objective while retaining format rewards to prevent forgetting issues. As discussed in Section 3.2.3, CAF establishes a trade-off between maximizing answer correctness and constraining retrieval counts. This incentivizes models to adaptively adjust retrieval depth based on query complexity, suppressing non-essential retrievals for simple queries while expanding information acquisition for complex multi-hop problems, leading to effective and efficient reasoning.
In summary, our phase-dependent training strategy resolves potential gradient conflicts, improving training stability and significantly boosting model performance. The pseudo-code of the training process is shown in Appendix B.1.
3.3. Hybrid Graph-Textual Retrieval
Compared to conventional RAG, GraphRAG exhibits superior relational modeling capabilities using graph structure. We integrate this advantage into both the training and generation phases of our framework. Specifically, we retrieve both structured triplets and relevant texts, replacing text fragments as commonly retrieved in existing RAGs and GraphRAGs with hybrid graph-textual retrieval.
By employing hybrid retrieval, our method has several advantages. Compared to lengthy text fragments, the compact nature of graph can drastically reduce token consumption, improving processing efficiency. This structured representation offers superior information density, allowing the model to focus directly on essential entity-relationships while inherently filtering out the redundant noise prevalent in raw text snippets. Therefore, the integration of structured knowledge significantly enhances the quality and efficiency of generated answers, as the reasoning operates on a cleaner and more relevant information foundation. We empirically verify the hybrid graph-textual retrieval in Section 4.3.3.
Table 1. The results of different methods for multi-hop question answering. The first three datasets are in-domain, and PopQA is out-of-domain (i.e., unseen during training). The best results are in bold and the second-best results are underlined.
Method HotpotQA MuSiQue 2Wiki PopQA
[…]
Vanilla LLM 3.08 30.00 44.32 2.01 15.50 43.14 4.55 29.50 44.37 10.46 25.90 50.07
Naive RAG 24.24 49.00 58.73 8.99 19.00 48.97 13.14 25.00 52.56 23.72 58.10 58.60
Prompt Engineering 22.22 41.00 58.67 7.72 25.50 49.91 17.54 36.00 54.44 13.37 27.50 51.61
SFT 12.10 49.50 51.76 8.29 21.00 48.90 11.40 48.50 49.46 10.45 54.10 50.56
KGP 10.73 21.00 50.92 4.61 11.00 46.70 10.16 18.00 50.27 21.01 50.00 56.10
ToG 11.44 21.50 50.48 5.02 8.00 47.05 14.45 21.50 53.11 29.21 47.00 62.94
LightRAG 7.96 18.50 49.15 3.09 6.00 44.99 7.58 10.00 47.95 14.55 34.30 53.07
PropRAG 25.47 55.00 59.76 10.11 21.50 50.77 17.17 35.00 53.85 23.29 57.30 58.22
G-retriever 14.81 21.50 54.89 3.38 5.00 47.47 15.12 19.00 52.77 2.14 6.80 44.63
HippoRAG2 27.52 53.00 61.02 8.90 23.50 50.19 12.32 29.00 51.92 23.99 57.46 58.16
R1-Searcher 26.82 35.00 59.15 12.35 14.50 50.32 15.89 22.50 51.51 23.24 43.20 56.80
GraphRAG-R1 38.00 52.00 66.66 20.06 26.50 57.49 32.24 49.50 63.56 35.04 60.00 65.31
4. Experiments
In this section, we conduct experiments to evaluate GraphRAG-R1, in order to answer the following research questions:
  • RQ1: Does GraphRAG-R1  effectively improve the performance of existing LLMs for both in-domain and out-of-domain datasets?
  • RQ2: Do all components, including PRA and CAF rewards, and phase-dependent training, contribute positively to our method?
  • RQ3: Is GraphRAG-R1  flexible enough to be seamlessly integrated with other retrievers?
4.1. Experimental Setup
4.1.1. Dataset
To comprehensively evaluate the GraphRAG-R1  framework, we follow (Jimenez Gutierrez et al., 2024; Wang, 2025) and select four widely used multi-hop QA benchmark datasets: HotpotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and PopQA (Mallen et al., 2023). For the training process, we utilize 80% of the training data from HotpotQA, 2Wiki, and MuSiQue. The remaining 20% of these datasets are employed as the in-domain testing set. Meanwhile, we utilize PopQA as an out-of-domain testing set, aiming to compare the generalization capabilities of different models. More details are provided in Appendix A.1.
4.1.2. Metrics
Following the common literature, we employ F1 Score (Wang et al., 2024), SBERT Similarity (SBERT) (Thakur et al., [n. d.]), and LLM-as-Judge Accuracy ($\mathrm{ACC_{L}}$)  (Song et al., 2025) as evaluation metrics. The F1 Score measures the lexical overlap between generated answers and ground-truth references. SBERT uses the cosine similarity between sentence embeddings generated by the SBERT model to measure semantic similarity. $\mathrm{ACC_{L}}$ leverages an LLM (specifically, Qwen-3-Turbo) as an evaluator to assess the correctness of predicted answers based on contextual understanding.
4.1.3. Baselines
We compare 7 representative GraphRAG methods:
  • KGP (Wang et al., 2024): It utilizes fine-tuned LLMs to explore neighboring nodes and aggregate information from a KG to produce answer.
  • ToG (Sun et al., [n. d.]): It dynamically expands and evaluates the most probable inference path in the KG to perform reasoning.
  • LightRAG (Guo et al., 2024): It enhances standard RAG by integrating graph-based text indexing with a dual-level retrieval strategy.
  • PropRAG (Wang, 2025): It optimizes the reasoning chain via propositional representation and beam search.
  • G-Retriever (He et al., 2024): It integrates LLM and graph neural network (GNN) with RAG for conversational graph QA.
  • HippoRAG2 (Gutiérrez et al., 2025): It uses personalized PageRank for deep passage integration inspired by the hippocampal indexing theory.
  • R1-Searcher (Song et al., 2025): It enhances the capability of LLMs to invoke RAG for retrieval through a two-stage RL framework.
We also compare with the following self-constructed baselines:
  • Vanilla LLM: It uses the raw backbone LLM for inference.
  • Naïve RAG: It adopts a single-hop retrieval from the vector database for generation.
  • Prompt Engineering: It adopts GraphRAG via structural prompts.
  • SFT: It adopts GraphRAG via SFT on the KGQA data.
4.1.4. Implementation Details
We utilize Qwen-2.5-7B as the default backbone LLM model. For baselines that require external closed-source LLMs, we utilize Qwen-3-Turbo. We use HippoRAG2 as the external retriever for LLM’s rollout retrieval.
[…]
4.2. Main Results
To answer Q1, we present the performance comparison between GraphRAG-R1  and baselines across four benchmarks in Table 1.We can draw the following three observations:
GraphRAG-R1  achieves consistent improvements across all datasets. Specifically, we achieve the best results for evaluation metrics F1 and SBERT on all datasets, and achieve three best results and one second-best result for $\mathrm{ACC_{L}}$. This fully validates the effectiveness of the GraphRAG-R1  framework in enhancing LLMs’ capability to handle multi-hop reasoning problems. Particularly, it outperforms the strongest baseline in F1 score by up to 38.03% on HotpotQA, 62.43% on MuSiQue, and 83.81% on 2Wiki, respectively. GraphRAG-R1  get more improvement on a more difficult multi-hop dataset. This demonstrates the framework’s efficacy in enhancing large language models’ multi-hop reasoning capabilities.
GraphRAG-R1 achieves leading performance on both in-domain datasets (HotpotQA, MuSiQue, and 2Wiki) and the out-of-domain dataset PopQA. Concretely, it outperforms the best baseline by up to 19.96% in terms of F1 on PopQA, indicating that our proposed method possesses strong generalization ability and can effectively handle unseen complex questions. In comparison, PropRAG, which show the most competitive performance on HotpotQA and MuSiQue, do not generalize to the out-of-domain dataset, indicating generalization is challenging for existing GraphRAG methods.
From the results of self-constructed baselines, existing techniques such as naive RAG, prompt engineering, and SFT can indeed improve the ability of vanilla LLM on complex question answering tasks. However, our method, which adopts RL-based training, is clearly more effective, outperforming all these non-RL approaches usually by a large margin. The consistent performance improvement validates that RL-based methods hold more potential in enhancing GraphRAG and the reasoning ability of LLMs in general.
Table 2. The ablation study results on PRA and CAF rewards. #Calls denotes the number of retrieval calls.
Method HotpotQA MuSiQue 2Wiki PopQA
[…]
GraphRAG-R1 38.00 1.94 20.06 2.36 32.24 2.05 35.04 1.49
- PRA 33.92 1.77 14.48 2.22 31.55 1.75 31.40 1.32
- CAF 28.36 2.33 19.82 2.80 22.16 2.36 26.98 2.05
- PRA & CAF 33.34 1.83 15.32 2.26 28.94 1.94 30.06 1.37
4.3. Further Analyses
To answer Q2, in this section, we conduct further analyses to verify the designs of our reward functions, phase-dependent training strategy, and hybrid retrieval. We also discuss the general compatibility of our method with various retrieval methods to answer Q3.
4.3.1. PRA & CAF Rewards
We perform ablation studies to evaluate the contribution of PRA and CAF reward designs. We remove PRA, CAF, and both rewards, denoted as “-PRA”, “-CAF”, and “-PRA & CAF”, respectively, and present the results in Table 2. The results show that removing either PRA or CAF, or both, leads to substantial performance drops in terms of F1 score. This confirms that both mechanisms are indispensable components of GraphRAG-R1 .
Table 3. The results of the ablation study on the phase-dependent training strategy.
Method HotpotQA MuSiQue 2Wiki PopQA
[…]
GraphRAG-R1 38.00 92.00 20.06 89.50 32.24 89.00 35.04 95.10
- cold start 30.31 89.00 14.84 89.50 30.04 89.50 28.83 92.80
- all stages 26.12 71.00 16.79 74.00 28.05 76.50 30.23 82.50
Besides the quantitative F1 score, we also delve deeper into the results by reporting the number of retrieval calls. We make the following observations. Without using PRA, our method has a reduced number of retrieval calls, indicating that it suffers from the “shallow retrieve” problem and leads to unsatisfactory performance. PRA dynamically regulates and encourages retrieval calls through its decay factor, making the retrieval behavior more controllable.
Table 4. The results of integrating GraphRAG-R1 with other GraphRAG retrieval methods. Note that GraphRAG-R1 is only trained once and do not access other RAG methods during training. Imp denotes the average improvement across all datasets.
Method Model HotpotQA MuSiQue 2Wiki PopQA Imp
[…]
Naive RAG Original 24.24 49.00 58.73 8.99 19.00 48.97 13.14 25.00 52.56 23.72 58.10 58.60
+GraphRAG-R1 32.08 48.00 62.72 20.77 28.00 56.59 30.14 37.50 61.22 31.96 57.50 52.96 + 18.04%
KGP Original 10.73 21.00 50.92 4.61 11.00 46.70 10.16 18.00 50.27 21.01 50.00 56.10
+GraphRAG-R1 36.16 41.50 65.81 13.33 14.50 53.82 27.79 35.50 60.41 18.33 55.00 55.39 + 36.25%
ToG Original 11.44 21.50 50.48 5.02 8.00 47.05 14.45 21.50 53.11 29.21 47.00 62.94
+GraphRAG-R1 16.82 25.50 55.36 8.29 8.50 51.77 21.95 26.00 57.68 22.30 30.50 60.10 + 3.52%
LightRAG Original 7.96 18.50 49.15 3.09 6.00 44.99 7.58 10.00 47.95 14.55 34.30 53.07
+GraphRAG-R1 20.85 27.00 57.40 8.59 12.50 48.95 19.64 23.00 55.31 28.42 49.00 60.50 + 38.37%
PropRAG Original 25.47 55.00 59.76 10.11 21.50 50.77 17.17 35.00 53.85 23.29 57.30 58.22
+GraphRAG-R1 30.87 42.00 62.25 18.78 21.00 55.32 34.86 45.50 64.44 32.69 57.80 63.44 + 13.16%
G-retriever Original 14.81 21.50 54.89 3.38 5.00 47.47 15.12 19.00 52.77 2.14 6.80 44.63
+GraphRAG-R1 13.52 15.50 54.64 7.49 2.50 51.56 15.23 17.50 54.07 5.86 11.10 49.06 + 3.66%
HippoRAG2 Original 27.52 53.00 61.02 8.90 23.50 50.19 12.32 29.00 51.92 23.99 57.46 58.16
+GraphRAG-R1 38.00 52.00 66.66 20.06 26.50 57.49 32.24 42.50 63.56 35.04 60.00 65.31 + 22.40%
On the other hand, without using CAF, the number of retrieval calls consistently increases across all datasets, indicating the model is more inclined to invoke external functions. However, this not yield to corresponding F1 improvements in answer quality. In fact, the F1 score is even lower than methods invoking few retrievals, indicating that the model tends towards “over-thinking” and performing unnecessary retrievals, potentially due to reward-hacking of solely using the PRA reward.
When both rewards are adopted, our method achieves the best performance with a suitable number of retrieval calls, indicating our method can calibrate its retrieval strategy to the inherent difficulty of the problem. For HotpotQA and PopQA, which contain relatively simpler problems, the number of retrieval calls is automatically decreased. Conversely, for inherently more complex problems demanding deeper information gathering, such as MuSiQue and 2Wiki, our model learns to invoke retrievals more frequently. In sum, our method can dynamically balance efficiency and effectiveness through reward designs.
4.3.2. Phase-Dependent Training Strategy
To validate the effectiveness of our training strategy, we compare with two variants: one without using the first cold-start stage (denoted as “-cold start”), and the other one without using stages, i.e., directly mixing all reward functions for training (denoted as “-all stages”). The results are shown in Table 3. We have the following observations.
Introducing the cold-start strategy effectively improves performance. In this stage, the model can learn simple tasks like format learning and retrieval invocation, which is beneficial for optimizing advanced policies in the later two stages.
Compared to training, which used format, PRA, and CAF rewards simultaneously, we found that the three-stage training strategy delivers superior performance. The first stage, which focused on format and tool invocation learning, establishes a more solid foundation for optimizing answer quality in the second stage. This approach avoids conflicts or convergence difficulties caused by overly complex reward signals.
4.3.3. Hybrid Graph-Textual Retrieval
To demonstrate the effectiveness of our hybrid graph-textual retrieval design, we compare different retrieval strategies. During training, we adopt three settings: using mixing text and graph (denoted as “Text+Graph”), text retrieval only (denoted as “Text”), and using graph only (denoted as ”Graph”). During inference, we utilize 4 different retrieval methods: using 5 text snippets (denoted as “Text5”), using 3 text snippets and 10 graph triplets (denoted as “Text3+Graph10”), using 1 text snippet and 10 graph triplets (denoted as “Text1+Graph10”), and using 20 graph triplets(denoted as “Graph20”). The results are shown in Figure 3 and we make the following two observations.
For different training strategies, training only with texts and training only with graph generally lead to performance decrease in the F1 score, demonstrating the effectiveness of mixed training .
For different inference strategies, as the proportion of graph data increases, the number of tokens consumed significantly reduces, while the F1 score of the model only slightly decreases. This demonstrates that graph information generally possess high information density, enabling the capture of valuable information needed for question answering. Our hybrid retrieval approach can achieve a suitable balance between model effectiveness and computational resource consumption.
Figure 3. A comparison of different retrieval formats. Left: the F1 score comparison. Right: the token cost for testing.
4.3.4. Integration with Other Retrievers
To answer Q3, we conduct experiments to integrate our proposed method with other existing retrieval methods. Specifically, we choose all RAG methods used in our baselines, including Naive RAG, ToG, KGP, LightRAG, PropRAG, G-retriever, and HippoRAG2. As shown in Table 4, GraphRAG-R1 generally improves the performance of different RAG methods across datasets. Specifically, GraphRAG-R1 achieves notable average improvements over the original methods, with the highest enhancement of 38.37% on LightRAG and consistent gains ranging from 3.52% to 36.25% in other baselines. Note that our method is only trained once and only has access to HippoRAG2 during training, indicating that the reasoning and retrieval ability learned by GraphRAG-R1 can be easily transferred and generalized to other GraphRAG methods.
More experiments including adopting LLaMA-3 as the backbone LLM, a comparison of the number of tokens, and detailed prompts used are provided in the Appendix.
5. Conclusion
In this paper, we propose GraphRAG-R1, an adaptive GraphRAG framework to enhance the multi-hop reasoning ability of LLMs through process-constrained RL. Our core designs include the rollout retrieval enhanced GRPO training method, process-constrained reward designs including PRA, CAF, and phase-depending training, and a hybrid graph-textual retrieval. Experimental results validate the superiority of GraphRAG-R1 over state-of-the-art baselines across both in-domain and out-of-domain datasets. We also demonstrate our method can be flexibly integrated with other retrieval methods. Future works include applying our method to more types of graphs and application scenarios.
Acknowledgements.
Appendix A More Experiments
A.1. More Details about Datasets
We adopt four widely used benchmark datasets, including HotpotQA, MuSiQue, and 2Wiki from HippoRAG (Jimenez Gutierrez et al., 2024) [^1] https://github.com/OSU-NLP-Group/HippoRAG/releases/tag/v1.0.0 and one simple QA dataset PopQA from PropRAG (Wang, 2025) [^2] https://github.com/ReLink-Inc/PropRAG/tree/main/outputs/popqa. Detailed dataset descriptions are as follows.
  • HotpotQA (Yang et al., 2018): A Wikipedia-based dataset where answering each question requires reasoning over multiple supporting documents.
  • MuSiQue (Trivedi et al., 2022): A challenging dataset consisting of questions composed of 2 to 4 single-hop questions, requiring multi-step reasoning.
  • 2Wiki (Ho et al., 2020): Another Wikipedia-based dataset where each question requires information from two articles to be correctly answered.
  • PopQA (Yang et al., 2018): A dataset of entity-centric QA pairs, where each question is generated by converting a knowledge tuple retrieved from Wikidata.
The statistics of these datasets are shown in Table 8.
A.2. Experiments with LLaMA-3
In this subsection, we investigate the effectiveness of GraphRAG-R1  across different base models. We adopt another representative open-sourced LLM, Llama-3-8B, as the backbone LLM. The experimental setting is similar to using Qwen-2.5-7B, except we add PRA reward in the third training stage to encourage the model to invoke retrieval, which enhances performance. The overall results are presented in Table 5 and the results of integrating with other retrievals are shown in Table 6. Collectively, these experiments demonstrate that our method is effective and outperforms baselines when using LLaMA-3-8B as the backbone LLM, showing that our method is generalizable across different LLM models.
Figure 4. System Prompt in GraphRAG-R1.
A.3. Cost and Efficiency Analysis
In this subsection, we evaluate the performance and cost of different reasoning-enhanced retrieval methods. We compare our approach with KGP and ToG, two reasoning-enhanced retrieval methods that also support multiple retrievals. In addition to comparing the performance measured by F1 scores, we also calculate the token consumption throughout the retrieval and reasoning processes. As shown in Table  7, GraphRAG-R1  achieves significant improvements in both performance and computational cost by deeply integrating the intrinsic reasoning capabilities of LLMs. Compared to baseline methods that also leverage LLM reasoning, our method demonstrates substantially improved answer quality while requiring lower token overhead than the other two methods. This validates the effectiveness and cost efficiency of our approach.
A.4. System Prompt in GraphRAG-R1
In this subsection we shown the detailed prompt of our model in Figure 4. This prompt is used to help the model gradually develop structured reasoning abilities through reinforcement learning. It encourages breaking down complex questions into subproblems and identifying what information is needed at each step. Through this design, the model learns to combine retrieval calls with internal reasoning abilities.
A.5. System Prompt for Generating Data with Desired Format
In this subsection, we introduce the prompt for generating data which is used in the first training stage to address the cold-start problem. Specifically, we utilize ERNIE-3.5-128K (Wang et al., 2021) model. The prompt design is shown in Figure 5. This structured prompt enables batch generation of training data. For each instance, we modify only the question and reference documents within the fixed template while retaining consistent few-shot examples. These examples explicitly define the required output format and logical structure. Following generation, we perform rigorous post-processing: validating output labels, correcting inconsistencies, and conducting quality assessment (e.g., relevance scoring).
Appendix B Details of GraphRAG-R1
Figure 5. System prompt for generating data with desired format used in the first training stage.
B.1. Algorithm Details
In this section, we provide the pseudo-code of our training pipeline in Algorithm 1.
Algorithm 1 GraphRAG-R1 Training Pipeline
1:Base LLM policy $\pi_{\text{ref}}$, Training Dataset $D_{\text{train}}$, External Retriever $R$, External LLM $\text{LLM}_{\text{gen}}$, Hyper-parameters: $\beta,R_{0},k,a,b$
2:Optimized LLM policy $\pi_{\theta}$
3:// Stage 1: Cold Start SFT (Format Following)
4:// Generate dataset $D_{\text{SFT}}$ with desired format:
5:for each $(q,\text{docs})\in D_{\text{train}}$ do
6:  Construct prompt with desired format (Fig 5)
7:  Generate $o_{\text{SFT}}$ with desired format using $\text{LLM}_{\text{gen}}$
8:  $D_{\text{SFT}}\leftarrow D_{\text{SFT}}\cup\{(q,o_{\text{SFT}})\}$
9:end for
10:Train $\pi_{\theta}$ with $D_{\text{SFT}}$ via supervised fine-tuning
11:// Stage 2: Behavior Shaping (PRA + Format Reward)
12:for each training epoch do
13:  for each $q\in D_{\text{train}}$ do
14:   $\text{SampleRollouts}(\pi_{\theta},q,R,G)$
15:   Calculate rewards using Eqs. (4)(5)
16:   Update policy via GRPO using Eqs. (2)(3)
17:  end for
18:end for
19:// Stage 3: Smartness Optimization (CAF + Format Reward)
20:for each training epoch do
21:  for each $q\in D_{\text{train}}$ do
22:   $\text{SampleRollouts}(\pi_{\theta},q,R,G)$
23:   Calculate rewards using Eqs. (4)(7)
24:   Update policy via GRPO Eqs. (2)(3)
25:  end for
26:end for
27:return $\pi_{\theta}$
28:// Function of Rollout with Thinking
[…]
30:  for $j=1$ to $G$ do
31:   $o_{j}\leftarrow[\ ]$
32:   while not terminated do
[…]
34:      $\text{query}\leftarrow\text{DecodeQuery}(o_{j})$
35:      $\text{docs}\leftarrow R(\text{query})$
[…]
37:     end if
38:     $o_{j}\leftarrow o_{j}\oplus\pi_{\theta}(\text{next\_token}|q,o_{j})$ // $\oplus$: concatenation
39:   end while
40:  end for
41:  return $\{o_{1},\dots,o_{G}\}$
42:end function

### v2

@@
 To handle the shallow retrieval problem, we design a Progressive Retrieval Attenuation (PRA) reward to encourage essential retrievals.
-Then, to handle the over-thinking problem, we design Cost-Aware F1 (CAF) reward to balance the model performance with computational costs.
+Then, to handle the over-thinking problem, we design a Cost-Aware F1 (CAF) reward to balance the model performance with computational costs.
 We further design a phase-dependent training strategy, containing three training stages corresponding to cold start and these two rewards.
@@
 Furthermore, our framework can be flexibly integrated with various existing retrieval methods, consistently delivering performance improvements.
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
+Reinforcement Learning, GraphRAG, Knowledge Graph
+journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792589isbn: 979-8-4007-2307-0/2026/04submissionid: rfp3241ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Machine learning
+Resource Availability: The source code has been made publicly available at https://doi.org/10.5281/zenodo.18349146 and can also be accessed via the project repository: https://github.com/ycygit/GraphRAG-R1.
+The model weights are available at https://doi.org/10.57967/hf/7629 and https://huggingface.co/yuchuanyue/GraphRAG-R1.
 1. Introduction
@@
 Despite remarkable progress, LLMs still exhibit significant limitations when faced with complex tasks that require integrating vast and diverse knowledge domains or require deep reasoning and multi-step logical inference (Talmor et al., 2019; Petroni et al., 2021).
-As shown in Figure 1(a), LLMs will provide an incorrect response when the question is over-complex and requires knowledge beyond their boundary.
+As shown in Figure 1(a), LLMs tend to provide an incorrect response when the question requires knowledge beyond their scope.
 Graph Retrieval-Augmented Generation (GraphRAG) (Xiang et al., 2025; Zhang et al., 2025a; Peng et al., 2024; Han et al., 2024; Zhu et al., 2024; Li et al., 2025), which utilizes knowledge graphs for knowledge representation, enables effective modeling of complex real-world relationships, such as causality, hierarchy, and dependency (Edge et al., 2024; Peng et al., 2024).
 However, existing GraphRAG methods still face significant performance bottlenecks when handling complex problems that require multi-hop reasoning.
-The main bottleneck is that existing GraphRAG methods rely on pre-defined or heuristic rules during query processing and graph retrieval, which are too rigid and inflexible.
+The main bottleneck is that existing methods rely on pre-defined or heuristic rules during query processing and retrieval, which are too rigid and inflexible.
 Concretely, during query processing, the semantic or lexical similarity search (Han et al., 2024; Gao et al., 2023b; Fan et al., 2024) remains the most widely adopted.
-For semantic similarity search, queries with high complexity are averaged during vectorization, making it difficult to gather accurate information for each distinct semantic element.
+For semantic similarity search, queries with high complexity are averaged out during vectorization, making it difficult to gather accurate information for each distinct semantic element.
 On the other hand, lexical similarity search captures only the superficial meaning of the query, easily losing implicit relationships and reasoning logic.
@@
 For instance, ToG (Sun et al., [n.
-d.]; Ma et al., 2024), a representative GraphRAG method, starts from a node in the knowledge graph and traverses the surrounding nodes based on breadth-first search strategy with pruning.
+d.]; Ma et al., 2024), a representative GraphRAG method, starts from a node in the graph and traverses the surrounding nodes based on breadth-first search strategy with pruning.
 While this strategy might work for short paths, it struggles with long-path problems, which require capturing logically sound but distant or even disconnected relationships.
@@
 To address this issue, we propose GraphRAG-R1, an adaptive GraphRAG framework with deep reasoning capabilities through process-constrained outcome-based RL.
-As shown in Figure 1(c), our method can decompose complex multi-hop problems by reasoning and autonomous, on-demand invocation of retrieval tools.
-Specifically, we employ a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking mechanism (Song et al., 2025) to autonomously leverage external retrieval tools on demand.
+As shown in Figure 1(c), our method can decompose complex multi-hop problems by reasoning and autonomous invocation of retrieval tools.
+Specifically, we employ a modified version of Group Relative Policy Optimization (GRPO) that supports rollout-with-thinking mechanism (Song et al., 2025) to leverage external retrieval tools on demand.
 Next, we design two process-constrained reward functions, Progressive Retrieval Attenuation (PRA) and Cost-Aware F1 (CAF), aiming to inject prior knowledge into the LLM’s training process.
 In detail, PRA aims to dynamically adjust the reward for retrieval actions.
-Initially, it encourages necessary retrievals to gather key information, preventing the model from shallow retrieval.
-As the reasoning deepens, PRA progressively attenuates retrieval rewards, guiding the model to rely more on acquired information for internal reasoning and avoid unnecessary retrieval overhead.
+It encourages necessary retrievals , preventing the model from shallow retrieval.
+As the reasoning deepens, PRA progressively attenuates retrieval rewards, guiding the model to avoid unnecessary retrieval overhead.
 On the other hand, CAF evaluates the final answer correctness using the F1 score while also accounting for the computational resources consumed during the reasoning and retrieval process.
 It encourages the model to achieve correct results more efficiently to prevent over-thinking.
-These two rewards collectively balance between effectiveness and computational cost.
+These two rewards collectively balance effectiveness and computational cost.
 We further design a three-stage phase-dependent training strategy.
 In the first stage, we focus on the answer format to alleviate cold start problems.
-In the second and third stage, we focus on improving the answer quality using the aforementioned two rewards.
-Lastly, our method employs a hybrid graph-textual retrieval instead of only retrieving textual information as in existing GraphRAGs.
+In the second and third stages, we focus on improving the answer quality using the aforementioned two rewards.
+Lastly, our method employs a hybrid graph-textual retrieval.
 By leveraging graph structures to precisely capture entity relationships while preserving the rich semantic details of the original text, our method provides more comprehensive and integrated knowledge for complex reasoning.
@@
   • We propose GraphRAG-R1, a pioneering framework that utilizes GRPO that supports rollout-with-thinking to enhance the reasoning ability of GraphRAG.
-To the best of our knowledge, we are the first to explore RL-based LLM training in GraphRAG.
   • We propose PRA and CAF, two tailored reward designs to guide the training process of RL and handle the shallow retrieval and over-thinking problems.
@@
 Figure 1.
-An example of comparing using only LLMs, GraphRAG, and GraphRAG-R1 in answering complex problems: (a) The LLM (Qwen-2.5-7b) directly produces an output, but the answer is incorrect, (b) HippoRAG2 (Gutiérrez et al., 2025) enhances the LLM (Qwen-2.5-7b) by retrieving external knowledge, but fail to generate correct outputs for complex problem, (c) our method successfully decomposes the problem and autonomously invokes retrieval methods, producing the correct output.
+An example of comparing using only LLMs, GraphRAG, and GraphRAG-R1 in answering complex problems: (a) The LLM (Qwen2.5-7B) directly produces an output, but the answer is incorrect, (b) HippoRAG2 (Gutiérrez et al., 2025) enhances the LLM (Qwen2.5-7B) by retrieving external knowledge, but fail to generate correct outputs for complex problem, (c) our method successfully decomposes the problem and autonomously invokes retrieval methods, producing the correct output.
+This figure provides a comparative schematic illustrating the performance of three distinct approaches in handling the same complex reasoning problem, visually demonstrating the superiority of the proposed GraphRAG-R1 framework.
+The schematic is divided into three juxtaposed sub-figures: (a) Using only the base Large Language Model (LLM, Qwen2.5-7B): The model directly generates an answer, which is incorrect.
+This reveals the limitations of LLMs regarding knowledge scope and multi-step reasoning.
+(b) Using an existing Graph Retrieval-Augmented Generation method (HippoRAG2) to augment the same LLM: While this method introduces external knowledge retrieval, it still fails to produce the correct answer for this complex problem.
+This highlights the bottlenecks of traditional GraphRAG methods in flexible problem decomposition and multi-hop reasoning.
+(c) Applying the proposed GraphRAG-R1 method: The model successfully decomposes the complex problem into sub-steps, autonomously and on-demand invokes external retrieval tools to acquire necessary information, and finally integrates the reasoning to produce the correct answer.
+Through this comparison, the figure clearly validates the effectiveness of GraphRAG-R1 in enhancing the LLM’s capability to solve complex problems.
 2. Related Work
@@
 For instance, GraphRAG based on knowledge graph (Pan et al., 2024) incorporates techniques like subgraph retrieval and Graph Neural Networks (GNNs) (He et al., 2024) to enhance the accuracy and explainability of LLMs in complex knowledge reasoning question answering.
-GraphRAG based on document graphs (Sonawane and Kulkarni, 2014) model semantic relationships between documents and internal document structures, effectively supporting long-document summarization and cross-document information integration (Sarthi et al., [n.
+GraphRAG based on document graphs (Sonawane and Kulkarni, 2014) models semantic relationships between documents and internal document structures, effectively supporting long-document summarization and cross-document information integration (Sarthi et al., [n.
 d.]).
-Furthermore, GraphRAG for specialized domains are increasingly trending.
-Examples include leveraging social graphs (Jiang and Ferrara, 2023; Zeng et al., 2024) to analyze user relationship for precision recommendations and constructing medical or disease graphs (Yang et al., 2024b; Li et al., 2024b) that integrate multi-source heterogeneous data to enable personalized treatment plan generation.
-However, existing GraphRAG methods still face significant performance bottlenecks when facing complex reasoning problem, primarily due to pre-defined or heuristic rules during query processing and graph retrieval that struggle with highly complex multi-hop problems.
+Furthermore, GraphRAG methods for specialized domains are increasingly trending.
+Examples include leveraging social graphs (Jiang and Ferrara, 2023; Zeng et al., 2024) to analyze user relationships for precision recommendations and constructing medical or disease graphs (Yang et al., 2024b; Li et al., 2024b) that integrate multi-source heterogeneous data to enable personalized treatment plan generation.
+However, existing GraphRAG methods still face significant performance bottlenecks when facing complex reasoning problems.
 For example, ToG (Sun et al., [n.
-d.]; Ma et al., 2024) introduces reasoning steps using breadth-first search (BFS) with pruning approach, but it heavily relies on prompt engineering and external LLMs and thus lacks flexibility.
+d.]; Ma et al., 2024) introduces reasoning steps using breadth-first search (BFS) with pruning approach, but it heavily relies on external LLMs and thus lacks flexibility.
 G-Retriever (He et al., 2024) adopts K-Nearest Neighbors-based retrieval with limited efficacy in multi-hop scenarios.
 Moreover, existing approaches suffer from shallow retrieval depth, since they often depend on predefined fixed paths or hop counts.
-HippoRAG (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025) optimizes knowledge integration through personalized PageRank but and only supports single-hop retrieval.
-KGP (Wang et al., 2024) and LightRAG (Guo et al., 2024) are typically confined to 2-3 hops, exhibiting significant performance degradation when handling problems that require deep-dependence mining.
+HippoRAG (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025) only supports single-hop retrieval.
+KGP (Wang et al., 2024) and LightRAG (Guo et al., 2024) are typically confined to 2-3 hops, exhibiting performance degradation when handling problems that require deep-dependence mining.
 In contrast, our work improves the reasoning ability of GraphRAG for complex problems using process-constrained RL, aiming to effectively resolves the issues of rigid retrieval strategies and complex reasoning shortcomings in existing GraphRAG approaches.
@@
 For example, DeepSeek-R1 (Guo et al., 2025) shows that large-scale RL can endow LLMs with deep thinking capabilities.
-Some pioneering work have also adopted RL to enhance RAG (Jin et al., 2025; Song et al., 2025).
+Some pioneering works have also adopted RL to enhance RAG (Jin et al., 2025; Song et al., 2025).
 For example, R1-Searcher (Song et al., 2025) and Search-R1 (Jin et al., 2025) achieve co-training of retrieval and generation, effectively improving the model’s ability to solve complex problems.
 However, a systematic exploration of RL schemes within the GraphRAG domain remains open.
-Technically speaking, directly extending existing RL methods into GraphRAG faces several challenges, Firstly, the existing methods usually employ a single outcome-oriented reward function, e.g., based solely on the correctness of final answer.
-However, this design becomes too simple for GraphRAG, since complex problem may involve multiple different retrieval calls to different components of the external knowledge graph.
-More critically, existing methods have often encountered with reward hacking in RL training, potentially resulting in shallow retrieval and over-thinking in GraphRAG.
-To tackle these issues, we propose GraphRAG-R1 with tailored reward functions and training strategy to non-trivially explore the potential of RL in enhancing GraphRAG, to the best of our knowledge, for the first time.
+Technically speaking, directly extending existing RL methods into GraphRAG faces several challenges.
+Firstly, the existing methods usually employ a single outcome-oriented reward function, e.g., based solely on the correctness of final answer.
+However, this design becomes too simple for GraphRAG, since a complex problem may involve complex retrieval calls of the external knowledge graph.
+More critically, existing methods have often encountered reward hacking in RL training, potentially resulting in shallow retrieval and over-thinking in GraphRAG.
+To tackle these issues, we propose GraphRAG-R1 with tailored reward functions and training strategy to non-trivially explore the potential of RL in enhancing GraphRAG.
 3. Method
@@
 An overview of GraphRAG-R1 : (a) the rollout retrieval enhanced GRPO as the training strategy of LLM, (b) process-constrained reward designs, containing PRA and CAF rewards and the phase-dependent training strategy, (c) the hybrid graph-textual retrieval, which is more informative than text fragments.
+This figure presents a comprehensive architectural overview of the proposed GraphRAG-R1 framework, illustrating its three core, interconnected components: (a) Rollout Retrieval Enhanced GRPO as the LLM Training Strategy: This component depicts the reinforcement learning (RL) training paradigm, specifically a modified Group Relative Policy Optimization (GRPO) algorithm.
+It visually represents how the LLM policy is trained via process-constrained outcome-based RL, enabling the model to autonomously invoke retrieval tools during its reasoning ”rollout” or thought process.
+(b) Process-Constrained Reward Designs: This section details the design of the reward functions that guide and constrain the model’s behavior during RL training.
+It highlights the two novel rewards and illustrates the phase-dependent training strategy that applies them sequentially (cold start, behavior shaping, smartness optimization) to effectively shape the model’s reasoning and retrieval patterns.
+(c) Hybrid Graph-Textual Retrieval Module: This part contrasts the proposed retrieval mechanism with traditional text-only retrieval.
+It visually demonstrates how combining structured graph triplets (capturing precise entity relationships) with unstructured text snippets (preserving rich semantic context) provides more comprehensive and integrated knowledge, thereby supplying higher-quality information for the LLM’s complex reasoning tasks.
 3.1. Rollout Retrieval Enhanced GRPO
@@
 More specifically, our method builds on the GRPO algorithm (Guo et al., 2025), which uses the average reward of multiple sampled outputs as a baseline.
-Specifically, for each input question $q$, GRPO samples a group of $G$ responses ${o_{1},o_{2},...,o_{G}}$ using the old policy $\pi_{\theta_{\text{old}}}$.
+During training, for each input query $q$, GRPO samples a group of $G$ responses $o_{1},o_{2},...,o_{G}$} using the old policy $\pi_{\theta_{\text{old}}}$.
 The policy model is then optimized by maximizing the following objective function:
@@
 $\text{clip}\left(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\right)\cdot A_{i}\bigg)-\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\bigg],$
-where the group size $G$ determines the number of responses generated per query, $\theta_{\text{old}}$ denotes the previous policy parameters, $\epsilon$ is the clipping range hyperparameter to constrain policy updates, and $A_{i}$ represents the advantage, computed based on the relative rewards of outputs within each group as follow:
+where the group size $G$ determines the number of responses generated per query, $\theta_{\text{old}}$ denotes the previous policy parameters, $\epsilon$ is the clipping range hyperparameter to constrain policy updates, and $A_{i}$ represents the advantage, computed based on the relative rewards of outputs within each group as follows:
 (3) $A_{i}=\frac{r_{i}-\mu_{r}}{\sigma_{r}},$
@@
 To better adapt GRPO to our GraphRAG scenario, we adopt two key modifications proposed by R1-Searcher (Song et al., 2025): Rollout-with-Thinking and Retrieval-Masked Loss.
-The core objective is to enable the model to autonomously leverage external retrieval tools on demand during inference, i.e., accessing supplementary knowledge when needed to solve problems beyond its internal capacity, and receive rewards for generating correct answers.
+The core objective is to enable the model to autonomously leverage external retrieval tools on demand during inference and receive rewards for generating correct answers.
 3.1.1. Rollout-with-Thinking
-The model is guided to utilize the external retrieval tool during generation by using two special tokens <begin_of_query>...<end_of_query> to indicate the call to the search tool.
+The model is guided to utilize the external retrieval tool during generation by using two special tokens <|begin_of_query|>...<|end_of_query|> to indicate the call to the search tool.
 When <|end_of_query|> appears, it signifies that the retrieval query has been fully generated.
@@
 This content then guides the model’s next step of reasoning.
-This method ensures retrieval is seamlessly integrated into the reasoning process, allowing the model to continue its reasoning based on the retrieved documents without interruption.
-Detailed prompts are provided in Appendix A.4
+This method ensures retrieval is seamlessly integrated into the reasoning process.
 3.1.2. Retrieval-Masked Loss
@@
 Applying the same optimization to retrieved tokens and LLM-generated tokens may negatively impact the training process.
-Therefore, <begin_of_documents> and <end_of_documents> are designated as special tokens and their enclosed content is masked during training, ensuring the policy gradient objective is calculated exclusively on LLM-generated tokens.
-This approach stabilizes training while preserving the flexibility of retrieval-augmented generation.
+Therefore, we utilize special tokens <|begin_of_documents|> and <|end_of_documents|> and their enclosed content is masked during training, ensuring the policy gradient objective is calculated only on LLM-generated tokens.
 3.2. Process-Constrained Reward Design
@@
 This reward function is designed to steer the model toward strictly adhering to the specified output format, ensuring that it learns to correctly invoke external retrieval and generate the final answer.
-The model should place the final response entirely within the <answer>...</answer> tags and guarantee that the entire output is fully readable and free of uninterpretable artifacts.
+The model should place the final response entirely within the <answer>...</answer> tags and ensure that the entire output is fully readable and free of uninterpretable artifacts.
 When invoking retrieval, the model should encapsulate retrieval queries within the <|begin_of_query|>...<|end_of_query|> tags with retrieved information contained within <|begin_of_documents|> ...
@@
 The model should explicitly invoke the retrieval mechanism through this query process rather than directly generating content without retrieval.
-An illustration of the format is shown in Figure 4 in Appendix A.4.
 The format reward is defined as follows:
@@
 (5) $R_{n}=\begin{cases}R_{0},&n=1\\ R_{n-1}+R_{0}\times k^{n-1},&n>1\end{cases},$
-[…]
+where $R_{0}$ is the base reward, $n$ is the number of retrieval calls, and $0<k<1$ is the decay factor.
+Through this dynamically adjusted reward weighting, the retrieval frequency of the model can be precisely controlled within the desired threshold ranges.
+Using the properties of geometric series, it is easy to see that:
 (6) $R_{0}\leq R_{n}<\frac{1}{1-k}R_{0},\;\forall n.$
@@
 The final reward is $R_{\text{retrieval}}=R_{N}$, where $N$ is the actual number of retrieval calls.
-Through this reward function, we can effectively restrict the number of model-triggered retrieval operations, effectively preventing shallow retrieval.
+Through this reward function, we can effectively restrict the number of retrieval calls, effectively preventing shallow retrieval.
 For different data characteristics, the hyperparameter $k$ can be flexibly adjusted to better steer the model’s learning process.
@@
 Although the PRA reward can regulate the retrieval frequency of the model within a reasonable range, this mechanism alone still relies on prior knowledge and could not fully activate the reasoning capabilities of LLMs.
-For example, the LLM may adopt unnecessarily numerous of retrievals, leading to the “over-thinking” problem and greatly waste the computational resources.
+For example, the LLM may adopt unnecessarily numerous retrievals, leading to the “over-thinking” problem and greatly waste the computational resources.
 An ideal GraphRAG model should deliver high-quality responses with precisely needed retrieval operations, which requires three core competencies: accurately assessing the problem difficulty, intelligently decomposing complex queries, and dynamically optimizing retrieval strategies.
@@
 where $F1$ is the F1 score of the generated answers, $a$ and $b$ are hyperparameters to control the coefficients in the regulatory function, and $N$ denotes the total number of retrieval operations.
-The CAF rewards ensure that responses attaining identical F1 scores receive higher rewards when using fewer retrievals, thereby incentivizing the model to proactively recognize problem complexity and refine retrieval decisions.
+The CAF reward ensures that responses achieving identical F1 score receive higher rewards when using fewer retrievals, thereby incentivizing the model to proactively recognize problem complexity and refine retrieval decisions.
 Consequently, CAF reward establishes a balance among reasoning depth, retrieval cost, and generation quality, effectively mitigating suboptimal behaviors of over-thinking.
@@
 In the second stage, we aim to shape the behavior of the model, synergistically combining the format reward and the PRA reward.
-This design recognizes standardized tool-calling protocols and retrieval behavior as essential pre-requisites for high-quality reasoning.
+This design recognizes standardized tool-calling protocols and retrieval behavior as essential prerequisites for high-quality reasoning.
 By learning with two objectives, the model can efficiently learn how and when to invoke the retrieval tools with structured output patterns.
@@
 In summary, our phase-dependent training strategy resolves potential gradient conflicts, improving training stability and significantly boosting model performance.
-The pseudo-code of the training process is shown in Appendix B.1.
 3.3. Hybrid Graph-Textual Retrieval
@@
 Compared to lengthy text fragments, the compact nature of graph can drastically reduce token consumption, improving processing efficiency.
-This structured representation offers superior information density, allowing the model to focus directly on essential entity-relationships while inherently filtering out the redundant noise prevalent in raw text snippets.
-Therefore, the integration of structured knowledge significantly enhances the quality and efficiency of generated answers, as the reasoning operates on a cleaner and more relevant information foundation.
+This structured representation allows the model to focus directly on essential entity-relationships while inherently filtering out the redundant noise.
+Therefore, the integration of structured knowledge significantly enhances the quality and efficiency of generated answers, as reasoning operates on a cleaner and more relevant information foundation.
 We empirically verify the hybrid graph-textual retrieval in Section 4.3.3.
@@
 Method HotpotQA MuSiQue 2Wiki PopQA
-[…]
+$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
 Vanilla LLM 3.08 30.00 44.32 2.01 15.50 43.14 4.55 29.50 44.37 10.46 25.90 50.07
@@
 SBERT uses the cosine similarity between sentence embeddings generated by the SBERT model to measure semantic similarity.
-$\mathrm{ACC_{L}}$ leverages an LLM (specifically, Qwen-3-Turbo) as an evaluator to assess the correctness of predicted answers based on contextual understanding.
+$\mathrm{ACC_{L}}$ leverages an LLM (specifically, Qwen3-Turbo) as an evaluator.
 4.1.3. Baselines
-We compare 7 representative GraphRAG methods:
-  • KGP (Wang et al., 2024): It utilizes fine-tuned LLMs to explore neighboring nodes and aggregate information from a KG to produce answer.
+We compare 7 representative methods:
+  • KGP (Wang et al., 2024): It utilizes fine-tuned LLMs to explore neighboring nodes and aggregate information from a KG to produce answers.
   • ToG (Sun et al., [n. d.]): It dynamically expands and evaluates the most probable inference path in the KG to perform reasoning.
@@
 4.1.4. Implementation Details
-We utilize Qwen-2.5-7B as the default backbone LLM model.
-For baselines that require external closed-source LLMs, we utilize Qwen-3-Turbo.
+We utilize Qwen2.5-7B as the default backbone LLM model.
+For baselines that require external closed-source LLMs, we use Qwen3-Turbo.
 We use HippoRAG2 as the external retriever for LLM’s rollout retrieval.
-[…]
+For training, in the first stage, we generate training data using ERNIE-3.5-128K (Wang et al., 2021).
+In the second stage, we set the base reward for PRA as $R_{0}=0.5$ and its decay factor as $k=0.5$.
+In the third stage, we set the hyper-parameter for CAF as $a=2$ and $b=0.1$.
+Throughout three stages, LoRA fine-tuning (Hu et al., 2021) is employed with $r=16$ and $\text{lora}\_{\text{alpha}}=32$.
+The sampling temperature is set to $1.0$ with a maximum retrieval count of $8$, the learning rate is 5e65e-6, and the number of epochs for SFT is $1$.
 4.2. Main Results
@@
 GraphRAG-R1  achieves consistent improvements across all datasets.
-Specifically, we achieve the best results for evaluation metrics F1 and SBERT on all datasets, and achieve three best results and one second-best result for $\mathrm{ACC_{L}}$.
+Specifically, we achieve the best results for evaluation metrics F1 and SBERT in all datasets, and achieve three best results and one third-best result for $\mathrm{ACC_{L}}$.
 This fully validates the effectiveness of the GraphRAG-R1  framework in enhancing LLMs’ capability to handle multi-hop reasoning problems.
-Particularly, it outperforms the strongest baseline in F1 score by up to 38.03% on HotpotQA, 62.43% on MuSiQue, and 83.81% on 2Wiki, respectively.
-GraphRAG-R1  get more improvement on a more difficult multi-hop dataset.
-This demonstrates the framework’s efficacy in enhancing large language models’ multi-hop reasoning capabilities.
-GraphRAG-R1 achieves leading performance on both in-domain datasets (HotpotQA, MuSiQue, and 2Wiki) and the out-of-domain dataset PopQA.
+Particularly, it outperforms the strongest baseline in F1 score by up to 38.08% on HotpotQA, 62.43% on MuSiQue, and 83.81% on 2Wiki, respectively.
+GraphRAG-R1  gets more improvement on a more difficult multi-hop dataset.
+GraphRAG-R1 achieves leading performance on both in-domain datasets and the out-of-domain dataset PopQA.
 Concretely, it outperforms the best baseline by up to 19.96% in terms of F1 on PopQA, indicating that our proposed method possesses strong generalization ability and can effectively handle unseen complex questions.
-In comparison, PropRAG, which show the most competitive performance on HotpotQA and MuSiQue, do not generalize to the out-of-domain dataset, indicating generalization is challenging for existing GraphRAG methods.
 From the results of self-constructed baselines, existing techniques such as naive RAG, prompt engineering, and SFT can indeed improve the ability of vanilla LLM on complex question answering tasks.
-However, our method, which adopts RL-based training, is clearly more effective, outperforming all these non-RL approaches usually by a large margin.
+However, our method, which adopts RL-based training, is clearly more effective, usually outperforming all these non-RL approaches by a large margin.
 The consistent performance improvement validates that RL-based methods hold more potential in enhancing GraphRAG and the reasoning ability of LLMs in general.
@@
 Method HotpotQA MuSiQue 2Wiki PopQA
-[…]
+$\mathrm{F1}$ $\mathrm{\#Calls}$ $\mathrm{F1}$ $\mathrm{\#Calls}$ $\mathrm{F1}$ $\mathrm{\#Calls}$ $\mathrm{F1}$ $\mathrm{\#Calls}$
 GraphRAG-R1 38.00 1.94 20.06 2.36 32.24 2.05 35.04 1.49
@@
 Method HotpotQA MuSiQue 2Wiki PopQA
-[…]
+$\mathrm{F1}$ $\mathrm{ACC_{F}}$ $\mathrm{F1}$ $\mathrm{ACC_{F}}$ $\mathrm{F1}$ $\mathrm{ACC_{F}}$ $\mathrm{F1}$ $\mathrm{ACC_{F}}$
 GraphRAG-R1 38.00 92.00 20.06 89.50 32.24 89.00 35.04 95.10
@@
 We make the following observations.
-Without using PRA, our method has a reduced number of retrieval calls, indicating that it suffers from the “shallow retrieve” problem and leads to unsatisfactory performance.
-PRA dynamically regulates and encourages retrieval calls through its decay factor, making the retrieval behavior more controllable.
+Without using PRA, our method has a reduced number of retrieval calls, indicating that it suffers from the “shallow retrieval” problem and leads to unsatisfactory performance.
+PRA makes the retrieval behavior more controllable.
 Table 4.
@@
 Method Model HotpotQA MuSiQue 2Wiki PopQA Imp
-[…]
+$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
 Naive RAG Original 24.24 49.00 58.73 8.99 19.00 48.97 13.14 25.00 52.56 23.72 58.10 58.60
@@
 ToG Original 11.44 21.50 50.48 5.02 8.00 47.05 14.45 21.50 53.11 29.21 47.00 62.94
-+GraphRAG-R1 16.82 25.50 55.36 8.29 8.50 51.77 21.95 26.00 57.68 22.30 30.50 60.10 + 3.52%
++GraphRAG-R1 16.82 25.50 55.36 8.29 8.50 51.77 21.95 26.00 57.68 22.30 37.00 60.10 + 5.26%
 LightRAG Original 7.96 18.50 49.15 3.09 6.00 44.99 7.58 10.00 47.95 14.55 34.30 53.07
@@
 G-retriever Original 14.81 21.50 54.89 3.38 5.00 47.47 15.12 19.00 52.77 2.14 6.80 44.63
-+GraphRAG-R1 13.52 15.50 54.64 7.49 2.50 51.56 15.23 17.50 54.07 5.86 11.10 49.06 + 3.66%
++GraphRAG-R1 13.52 15.50 54.64 10.56 3.50 54.73 15.23 17.50 54.07 5.86 11.10 49.06 + 6.18%
 HippoRAG2 Original 27.52 53.00 61.02 8.90 23.50 50.19 12.32 29.00 51.92 23.99 57.46 58.16
@@
 However, this not yield to corresponding F1 improvements in answer quality.
-In fact, the F1 score is even lower than methods invoking few retrievals, indicating that the model tends towards “over-thinking” and performing unnecessary retrievals, potentially due to reward-hacking of solely using the PRA reward.
+In fact, the F1 score is even lower than methods invoking few retrievals, indicating that the model tends towards “over-thinking” and performing unnecessary retrievals.
 When both rewards are adopted, our method achieves the best performance with a suitable number of retrieval calls, indicating our method can calibrate its retrieval strategy to the inherent difficulty of the problem.
@@
 Conversely, for inherently more complex problems demanding deeper information gathering, such as MuSiQue and 2Wiki, our model learns to invoke retrievals more frequently.
-In sum, our method can dynamically balance efficiency and effectiveness through reward designs.
 4.3.2. Phase-Dependent Training Strategy
-To validate the effectiveness of our training strategy, we compare with two variants: one without using the first cold-start stage (denoted as “-cold start”), and the other one without using stages, i.e., directly mixing all reward functions for training (denoted as “-all stages”).
-The results are shown in Table 3.
-We have the following observations.
+To validate the effectiveness of our training strategy, we compare our method with two variants: one without the first cold-start stage (denoted as “-cold start”), and the other without distinct stages, i.e., directly mixing all reward functions for training (denoted as “-all stages”).
+The results are shown in Table  3.
+Here, $\mathrm{ACC_{F}}$ represents the percentage of model outputs that are entirely accurate in format.
 Introducing the cold-start strategy effectively improves performance.
@@
 Compared to training, which used format, PRA, and CAF rewards simultaneously, we found that the three-stage training strategy delivers superior performance.
-The first stage, which focused on format and tool invocation learning, establishes a more solid foundation for optimizing answer quality in the second stage.
 This approach avoids conflicts or convergence difficulties caused by overly complex reward signals.
@@
 To demonstrate the effectiveness of our hybrid graph-textual retrieval design, we compare different retrieval strategies.
-During training, we adopt three settings: using mixing text and graph (denoted as “Text+Graph”), text retrieval only (denoted as “Text”), and using graph only (denoted as ”Graph”).
-During inference, we utilize 4 different retrieval methods: using 5 text snippets (denoted as “Text5”), using 3 text snippets and 10 graph triplets (denoted as “Text3+Graph10”), using 1 text snippet and 10 graph triplets (denoted as “Text1+Graph10”), and using 20 graph triplets(denoted as “Graph20”).
-The results are shown in Figure 3 and we make the following two observations.
+During training, we adopt three settings: using mixing text and graph (“Text+Graph”), text retrieval only (“Text”), and using graph only (”Graph”).
+During inference, we utilize 4 different retrieval methods: using 5 text snippets (“Text5”), using 3 text snippets and 10 graph triplets (“Text3+Graph10”), using 1 text snippet and 10 graph triplets (“Text1+Graph10”), and using 20 graph triplets ( “Graph20”).
+The results are shown in Figure  3.
 For different training strategies, training only with texts and training only with graph generally lead to performance decrease in the F1 score, demonstrating the effectiveness of mixed training.
 For different inference strategies, as the proportion of graph data increases, the number of tokens consumed significantly reduces, while the F1 score of the model only slightly decreases.
-This demonstrates that graph information generally possess high information density, enabling the capture of valuable information needed for question answering.
-Our hybrid retrieval approach can achieve a suitable balance between model effectiveness and computational resource consumption.
+This demonstrates that our hybrid retrieval approach can achieve a suitable balance between model effectiveness and computational resource consumption.
 Figure 3. A comparison of different retrieval formats. Left: the F1 score comparison. Right: the token cost for testing.
+[…]
 4.3.4. Integration with Other Retrievers
 To answer Q3, we conduct experiments to integrate our proposed method with other existing retrieval methods.
-Specifically, we choose all RAG methods used in our baselines, including Naive RAG, ToG, KGP, LightRAG, PropRAG, G-retriever, and HippoRAG2.
 As shown in Table 4, GraphRAG-R1 generally improves the performance of different RAG methods across datasets.
-Specifically, GraphRAG-R1 achieves notable average improvements over the original methods, with the highest enhancement of 38.37% on LightRAG and consistent gains ranging from 3.52% to 36.25% in other baselines.
-Note that our method is only trained once and only has access to HippoRAG2 during training, indicating that the reasoning and retrieval ability learned by GraphRAG-R1 can be easily transferred and generalized to other GraphRAG methods.
-More experiments including adopting LLaMA-3 as the backbone LLM, a comparison of the number of tokens, and detailed prompts used are provided in the Appendix.
+The last column ”Imp” is defined as the average percentage increase in all metrics after integrating GraphRAG-R1  relative to original method.
+Specifically, GraphRAG-R1 achieves notable average improvements over the original methods, with the highest improvement of 38.37% on LightRAG and consistent gains ranging from 5.26% to 36.25% in other baselines.
+Note that our method is only trained once and only has access to HippoRAG2 during training, indicating that the reasoning and retrieval ability learned by GraphRAG-R1 can be easily generalized to other GraphRAG methods.
+More experiments are provided in the Appendix.
 5. Conclusion
@@
 Acknowledgements.
+This work was supported in part by the National Natural Science Foundation of China (No.
+92570113, 62472018), the Natural Science Foundation of Tianjin (No.
+24JCQNJC02170), and the Tianjin Science and Technology Plan Project (No.
+24HHXCSS00001).
 Appendix A More Experiments
-A.1. More Details about Datasets
+A.1. Experiments with Qwen2.5-7B-Instruct
+In this study, we further explore the performance of GraphRAG-R1 on instruct-tuned models.
+We observed that results based on the original Qwen2.5-7B-base still have room for improvement on certain tasks, particularly in terms of instruction compliance and contextual understanding.
+Therefore, we selected the Qwen2.5-7B-Instruct model as the backbone.
+Compared to the base model, the instruct-tuned model demonstrates stronger capabilities in understanding user intent, following output formats, and handling multi-turn dialogues.
+Based on this model, we retained the original GraphRAG-R1 framework and only performed adaptive tuning on the training strategy, prompt templates, and inference hyperparameters, aiming to verify the stability and scalability of our method across different types of backbone models.
+The experimental results demonstrate that the tuned GraphRAG-R1 with Qwen2.5-7B-Instruct not only significantly outperforms the base version but also achieves the best performance on most evaluation metrics across datasets such as HotpotQA, MuSiQue, 2Wiki, and PopQA.
+Detailed performance comparisons are provided in Table 5.
+Furthermore, we tested the integration of our approach with other retrievers, as shown in Table 6.
+In all tested configurations, GraphRAG-R1 consistently improved performance over the original retrieval methods, demonstrating strong transferability and generalization.
+This indicates that our framework can be flexibly adapted to different retrieval mechanisms.
+A.2. Experiments with LLaMA-3-8B
+Table 7. The results of different methods for multi-hop question answering using Llama-3-8B as the backbone LLM. The best results are in bold and the second-best results are underlined.
+Method HotpotQA MuSiQue 2Wiki PopQA
+$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
+Naive RAG 9.57 27.00 48.25 1.17 4.00 40.11 4.14 10.00 43.58 8.14 24.70 46.59
+KGP 3.61 17.00 46.41 1.38 3.50 44.49 8.23 26.50 48.48 10.75 59.00 48.13
+ToG 8.97 16.50 49.83 3.28 10.00 44.74 7.91 23.00 48.36 10.96 37.50 49.57
+LightRAG 5.33 12.50 43.95 1.83 3.50 41.51 4.68 12.00 43.95 6.71 27.50 45.83
+PropRAG 12.47 23.00 49.03 3.83 6.50 42.17 5.06 11.50 43.06 9.76 25.70 48.10
+G-retriever 1.91 23.00 43.09 1.06 12.50 42.77 2.43 18.00 42.26 0.25 38.10 39.32
+HippoRAG2 6.89 22.00 46.06 1.85 4.00 42.03 5.61 12.00 43.51 7.85 26.10 46.57
+GraphRAG-R1 13.16 26.50 53.06 6.32 16.00 48.96 13.36 25.00 51.46 11.64 41.50 50.42
+In this subsection, we investigate the effectiveness of GraphRAG-R1  across different base models.
+We adopt another representative open-sourced LLM, Llama-3-8B, as the backbone LLM.
+The experimental setting is similar to using Qwen2.5-7B, except we add PRA reward in the third training stage to encourage the model to invoke retrieval, which enhances performance.
+The overall results are presented in Table 7.
+Collectively, these experiments demonstrate that our method is effective and outperforms baselines when using LLaMA-3-8B as the backbone LLM, showing that our method is generalizable across different LLM models.
+A.3. Hyperparameter Sensitivity Analysis
+To evaluate the hyperparameter robustness of the GraphRAG-R1 framework, we conducted sensitivity analysis on four key parameters.
+The results are shown in Figure 4 and demonstrate that the performance of GraphRAG-R1 has a wide robustness interval around the parameters selected in the main experiment, which enhances the reproducibility and practical reliability of the method.
+Figure 4.
+Sensitivity analysis of key hyperparameters in GraphRAG-R1.
+Subfigures (a)-(d) illustrate the variation in F1 score on the four validation sets as the key hyperparameters are varied within specified ranges.
+Other hyperparameters were fixed at their main experiment values.
+The gray dashed lines indicate the parameter values selected for the main experiments.
+Sensitivity analysis of key hyperparameters in GraphRAG-R1.
+Subfigures (a)-(d) illustrate the variation in F1 score on the four validation sets as the key hyperparameters are varied within specified ranges.
+Other hyperparameters were fixed at their main experiment values.
+The gray dashed lines indicate the parameter values selected for the main experiments.
+A.4. Cost and Efficiency Analysis
+Table 8. Performance and cost comparison with KGP and ToG, two reasoning-enhanced GraphRAG methods
+Method HotpotQA MuSiQue 2Wiki PopQA
+$\mathrm{F1}$ $\mathrm{Token}$ $\mathrm{F1}$ $\mathrm{Token}$ $\mathrm{F1}$ $\mathrm{Token}$ $\mathrm{F1}$ $\mathrm{Token}$
+KGP 10.73 2687.44 4.61 2858.23 10.16 2173.02 21.01 3182.02
+ToG 11.44 3136.61 5.02 2873.56 14.45 2065.94 29.21 2838.49
+GraphRAG-R1 38.00 1674.47 20.06 2086.10 32.24 1883.21 35.04 1449.80
+Table 9.
+The results of integrating GraphRAG-R1 with other GraphRAG retrieval methods.
+Note that GraphRAG-R1 is only trained once and do not access other RAG methods during training.
+Imp denotes the average improvement across all datasets.
+Method Model HotpotQA MuSiQue 2Wiki PopQA Imp
+$\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$ $\mathrm{F1}$ $\mathrm{ACC_{L}}$ $\mathrm{SBERT}$
+Dalk Original 0.62 7.50 41.96 0.40 3.00 40.64 0.75 4.00 41.66 0.14 8.80 39.37
++GraphRAG-R1 15.95 20.50 52.38 8.15 8.50 48.33 10.45 14.40 49.54 13.38 26.60 51.90 + 69.50%
+In this subsection, we evaluate the performance and cost of different reasoning-enhanced retrieval methods.
+We compare our approach with KGP and ToG, two reasoning-enhanced retrieval methods that also support multiple retrievals.
+In addition to comparing the performance measured by F1 scores, we also calculate the token consumption throughout the retrieval and reasoning processes.
+Specifically, for ToG and KGP, we count the number of accesses to the knowledge graph, for our method, we tally the frequency of calls to external retrieval tools.
+Token overhead accounts for consumption throughout the retrieval and reasoning processes.
+As shown in Table  8, GraphRAG-R1  achieves significant improvements in both performance and computational cost by deeply integrating the intrinsic reasoning capabilities of LLMs.
+Compared to baseline methods that also leverage LLM reasoning, our method demonstrates substantially improved answer quality while requiring lower token overhead than the other two methods.
+This validates the effectiveness and cost efficiency of our approach.
+A.5. Cross-Domain Adaptation Verification
+GraphRAG-R1 is designed as a general framework agnostic to specific graph structures, tasks, and domains.
+To verify its domain generalization and compatibility, we adapted an advanced retrieval method DALK (Li et al., 2024b) originally designed for the medical domain, integrated it into our framework, and evaluated it on general-domain datasets.
+We only replacing GraphRAG-R1’s underlying retrieval tool with the adapted DALK retriever.
+As shown in Table 9, the integrated framework achieved a significant improvement compared to the original DALK method, with a relative F1 score increase of approximately 69.50% on Qwen2.5-7B.
+This strongly proves that our framework can flexibly compatibly integrate and enhance specialized retrievers from different domains.
+A.6. More Details about Datasets
+Table 10. Dataset Statistics
+Metric MuSiQue 2Wiki HotpotQA PopQA
+Graph # Passages 11,656 6,119 9,221 8,676
+# Nodes 91,729 42,694 82,157 74,472
+# Triplets 107,448 50,671 98,709 139,273
+Questions Total 1,000 1,000 1,000 1,000
+Training Set 800 800 800 0
+Testing Set 200 200 200 1,000
 We adopt four widely used benchmark datasets, including HotpotQA, MuSiQue, and 2Wiki from HippoRAG (Jimenez Gutierrez et al., 2024) [^1] https://github.com/OSU-NLP-Group/HippoRAG/releases/tag/v1.0.0 and one simple QA dataset PopQA from PropRAG (Wang, 2025) [^2] https://github.com/ReLink-Inc/PropRAG/tree/main/outputs/popqa.
-Detailed dataset descriptions are as follows.
+Detailed descriptions of the datasets are as follows.
   • HotpotQA (Yang et al., 2018): A Wikipedia-based dataset where answering each question requires reasoning over multiple supporting documents.
@@
   • 2Wiki (Ho et al., 2020): Another Wikipedia-based dataset where each question requires information from two articles to be correctly answered.
-  • PopQA (Yang et al., 2018): A dataset of entity-centric QA pairs, where each question is generated by converting a knowledge tuple retrieved from Wikidata.
+  • PopQA (Mallen et al., 2023): A dataset of entity-centric QA pairs, where each question is generated by converting a knowledge tuple retrieved from Wikidata.
 The statistics of these datasets are shown in Table 10.
-A.2. Experiments with LLaMA-3
-In this subsection, we investigate the effectiveness of GraphRAG-R1  across different base models.
-We adopt another representative open-sourced LLM, Llama-3-8B, as the backbone LLM.
-The experimental setting is similar to using Qwen-2.5-7B, except we add PRA reward in the third training stage to encourage the model to invoke retrieval, which enhances performance.
-The overall results are presented in Table 5 and the results of integrating with other retrievals are shown in Table 6.
-Collectively, these experiments demonstrate that our method is effective and outperforms baselines when using LLaMA-3-8B as the backbone LLM, showing that our method is generalizable across different LLM models.
-Figure 4. System Prompt in GraphRAG-R1.
-A.3. Cost and Efficiency Analysis
-In this subsection, we evaluate the performance and cost of different reasoning-enhanced retrieval methods.
-We compare our approach with KGP and ToG, two reasoning-enhanced retrieval methods that also support multiple retrievals.
-In addition to comparing the performance measured by F1 scores, we also calculate the token consumption throughout the retrieval and reasoning processes.
-As shown in Table  7, GraphRAG-R1  achieves significant improvements in both performance and computational cost by deeply integrating the intrinsic reasoning capabilities of LLMs.
-Compared to baseline methods that also leverage LLM reasoning, our method demonstrates substantially improved answer quality while requiring lower token overhead than the other two methods.
-This validates the effectiveness and cost efficiency of our approach.
-A.4. System Prompt in GraphRAG-R1
-In this subsection we shown the detailed prompt of our model in Figure 4.
-This prompt is used to help the model gradually develop structured reasoning abilities through reinforcement learning.
-It encourages breaking down complex questions into subproblems and identifying what information is needed at each step.
-Through this design, the model learns to combine retrieval calls with internal reasoning abilities.
-A.5. System Prompt for Generating Data with Desired Format
-In this subsection, we introduce the prompt for generating data which is used in the first training stage to address the cold-start problem.
-Specifically, we utilize ERNIE-3.5-128K (Wang et al., 2021) model.
-The prompt design is shown in Figure 5.
-This structured prompt enables batch generation of training data.
-For each instance, we modify only the question and reference documents within the fixed template while retaining consistent few-shot examples.
-These examples explicitly define the required output format and logical structure.
-Following generation, we perform rigorous post-processing: validating output labels, correcting inconsistencies, and conducting quality assessment (e.g., relevance scoring).
-Appendix B Details of GraphRAG-R1
-Figure 5. System prompt for generating data with desired format used in the first training stage.
-B.1. Algorithm Details
-In this section, we provide the pseudo-code of our training pipeline in Algorithm 1.
-Algorithm 1 GraphRAG-R1 Training Pipeline
-1:Base LLM policy $\pi_{\text{ref}}$, Training Dataset $D_{\text{train}}$, External Retriever $R$, External LLM $\text{LLM}_{\text{gen}}$, Hyper-parameters: $\beta,R_{0},k,a,b$
-2:Optimized LLM policy $\pi_{\theta}$
-3:// Stage 1: Cold Start SFT (Format Following)
-4:// Generate dataset $D_{\text{SFT}}$ with desired format:
-5:for each $(q,\text{docs})\in D_{\text{train}}$ do
-6:  Construct prompt with desired format (Fig 5)
-7:  Generate $o_{\text{SFT}}$ with desired format using $\text{LLM}_{\text{gen}}$
-8:  $D_{\text{SFT}}\leftarrow D_{\text{SFT}}\cup\{(q,o_{\text{SFT}})\}$
-9:end for
-10:Train $\pi_{\theta}$ with $D_{\text{SFT}}$ via supervised fine-tuning
-11:// Stage 2: Behavior Shaping (PRA + Format Reward)
-12:for each training epoch do
-13:  for each $q\in D_{\text{train}}$ do
-14:   $\text{SampleRollouts}(\pi_{\theta},q,R,G)$
-15:   Calculate rewards using Eqs. (4)(5)
-16:   Update policy via GRPO using Eqs. (2)(3)
-17:  end for
-18:end for
-19:// Stage 3: Smartness Optimization (CAF + Format Reward)
-20:for each training epoch do
-21:  for each $q\in D_{\text{train}}$ do
-22:   $\text{SampleRollouts}(\pi_{\theta},q,R,G)$
-23:   Calculate rewards using Eqs. (4)(7)
-24:   Update policy via GRPO Eqs. (2)(3)
-25:  end for
-26:end for
-27:return $\pi_{\theta}$
-28:// Function of Rollout with Thinking
-[…]
-30:  for $j=1$ to $G$ do
-31:   $o_{j}\leftarrow[\ ]$
-32:   while not terminated do
-[…]
-34:      $\text{query}\leftarrow\text{DecodeQuery}(o_{j})$
-35:      $\text{docs}\leftarrow R(\text{query})$
-[…]
-37:     end if
-38:     $o_{j}\leftarrow o_{j}\oplus\pi_{\theta}(\text{next\_token}|q,o_{j})$ // $\oplus$: concatenation
-39:   end while
-40:  end for
-41:  return $\{o_{1},\dots,o_{G}\}$
-42:end function